OpenStudy (anonymous):

If (a+3) divides evenly into ka^3−10a^2+2a+3 , find k

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

heeeeeeeeelp please @Loser66

OpenStudy (loser66):

let it =0, replace a = -3 to find k

OpenStudy (anonymous):

k=10.3 ?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

well is it??

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

I got k =-93/27

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

I got +93/27

OpenStudy (anonymous):

not (-)

OpenStudy (loser66):

ka^3-10a^2+2a+3 =0 ka^3 = 10a^2-2a-3 k = (10a^2-2a-3)/ a^3 replace a = -3 k = (10*(-3)^2-2(-3)-3)/ (-3)^3= 93/ (-27) = -93/27

OpenStudy (loser66):

Yes / No??

OpenStudy (anonymous):

yes

OpenStudy (loser66):

ok

OpenStudy (anonymous):

-93/27 = -31/9 Loser66 posted first. Refer to the attachment.